A conducting loop is made in the form of two squares of sides s1 = 3.8cm and s2 = 7.9 cm as shown. At time t = 0, the loop enters a region of length L = 20.5 cm that contains a uniform magnetic field B = 1.1 T, directed in the positive z-direction. The loop continues through the region with constant speed v = 37 cm/s. The resistance of the loop is R = 1.1 Ω.

h16_constantB

A.) At time t = t1 = 0.034 s, what is I1, the induced current in the loop? I1 is defined to be positive if it is in the counterclockwise direction.

B.) At time t = t2 = 0.728 s, what is I2, the induced current in the loop? I2 is defined to be positive if it is in the counterclockwise direction.

C.) What is Fx(t2), the x-component of the force that must be applied to the loop to maintain its constant velocity v = 37 cm/s at t = t2 = 0.728 s?

D.) At time t = t3 = 0.588 s, what is I3, the induced current in the loop? I3 is defined to be positive if it is in the counterclockwise direction.
Answer #1

A) At t = t1 = 0.034 s, small square will be inside the magnetic field.

induced emf in the loop, emf = B*v*s1


induced current, I1 = emf/R

= B*v*s1/R

= 1.1*0.37*3.8*10^-2/1.1

= 0.014 A

direction : clockwise

so, I1 = -0.014 A

B)

At t = t2 = 0.718 s, big square will be inside the magnetic field and moving out.

induced emf in the loop, emf = B*v*s2


induced current, I2 = emf/R

= B*v*s2/R

= 1.1*0.37*7.9*10^-2/1.1

= 0.0292 A

direction : counter clockwise

so, I2 = +0.0292 A

C) F(x) = B*I2*s2*sin(90)

= 1.1*0.0292*7.9*10^-2*1

= 2.53*10^-3 N

D)

At t = t3 = 0.588 s, small square will be inside the magnetic field and coming out.

induced emf in the loop, emf = B*v*s1

induced current, I3 = emf/R

= B*v*s1/R

= 1.1*0.37*3.8*10^-2/1.1

= 0.014 A

direction : counter clockwise

so, I3 = +0.014 A